You are CEO of Eastco, and you recently paid $58,000 or about 2x revenue (well under industry average) to purchase Westco, which makes a product nearly identical to yours. This move will allow you to operate in both regions (East and West) of the country.

If you combine the markets, the total Fixed Cost would be $11,000 per month, and your larger purchase size will allow you to advantageous terms for your raw materials, meaning VC will be $4 per unit.

Demand and MR in the combined market is as follows:
Qd = 600 - 6P
MR = 100 - 0.333Q

a) Compute your profit in the combined market.

Additional analysis shows that demand and fixed costs are different in the 2 regions.

The West has the following demand and Mg:
Qd = 300 - 4P
MR = 75 - 0.5Q

Fixed costs associated with operating in the West are $5000/month.

While the East is a less price sensitive market:
Qd = 300 - 2P
MR = 150 - Q

Fixed costs associated with operating in the East are $6000/month.

The area manager from the East suggests you use a strategy that charges each market a separate price. If you operate in separate markets, the VC associated with each unit is $5.

b) Calculate your profits using this strategy. What is this strategy called?

c) What strategy do you suggest in the Long Run?
Answer #1

A).

The MR under the combine market is given by, => MR = 100 – 0.333*q and the AVC = 4, => TVC = 4*q, => MC = 4. So, at the equilibrium “MR=MC”.

=> 100 – 0.333*q = 4, => 0.333*q = 96, => q = 96/0.333 = 288.29, => q = 288.29. the profit maximizing price is given by, => P = 100-q/6 = 100-288.29/6 = 51.95.

=> the profit is given by “A = Pq – C = 51.95*288.29 – [11,000 + 4*288.29] = 14,976.66 – [12,153.16] = 2,823.5 > 0. So, the profit under the combined market is given by “A = 2,823.5”.

B).

The MR and MC is “west” are given by, “MR=75-0.5*q” and “MC=5”, => at the equilibrium “MR=MC”.

=> 75 - q/2=5, => q/2 = 70, => q=140. Now, the demand function is given by “P = 75 - q/4 = 75 – 140/4 = 40. So, the optimum price and the quantity are given by “P=40” and “q=140” in “WEST”.

Now, the MR and MC is “east” are given by, “MR = 150 - q” and “MC=5”, => at the equilibrium “MR=MC”.

=> 150 – q = 5, => q = 145. Now, the demand function is given by “P = 150 - q/2 = 150 – 145/2 = 77.5. So, the optimum price and the quantity are given by “P=77.5” and “q=145” in “EAST”.

=> the profit is given by.

=> A = TR – TC = 40*140 + 77.5*145 – [5000+6000+5*(140+145)] = 16,837.5 – [11,000+1,425] = 4,412.5 > 0.

C).

Here the profit under the combined market is less than separate market, => charging separate price to separate market will be much profitable in the LR.